I need help on solving linear equations with abslute value's in them the problem is 3[y+51]-15=6 []used to represent abslute value brackets. I dont need the answer I just need help
Answer by khwang(438)

First of all, convert 3|y+51|-15=6 to
3|y+51| = 21 by adding 15 on both sides
then dividing by 7:
|y + 51| = 7
By definition of absolute value,we see that |x| = b is equivalent
to the equations x = b or x = -b.
Hence,|y + 51| = 7 is equivalent
to y + 51 = 7...(1)
or y + 51 = -7...(2)
From (1): y = -44
From (1): y = -58